OpenStudy (anonymous):

A. A line with a slope of -2/3 and a y-intercept of 4; in slope-intercept form.... can anyone help??:/

OpenStudy (anonymous):

y=mx+b y=-2/3x+4

pooja195 (pooja195):

y=mx+b m=slope b=y intercept